National Advertising just paid a dividend of D0 = $0.75 per share, and that dividend is expected to grow at a constant rate of 6.50% per year in the future. The company's beta is 0.8, the required return on the market is 10.50%, and the risk-free rate is 4.50%. What is the company's current stock price? Select the correct answer. a. $29.21 b. $28.53 c. $28.87 d. $29.55 e. $28.19
Answer #1

required return= risk-free rate +Beta*(market rate- risk-free rate )

=4.5+0.8*(10.5-4.5)

=9.3%

Current price=D1/(Required return-Growth rate)

=(0.75*1.065)/(0.093-0.065)

=$28.53(Approx).